# On the adversarial robustness of Locality-Sensitive Hashing in Hamming space

### v2

@@ Abstract
Locality-sensitive hashing [IM98] is a classical data structure for approximate nearest neighbor search. It allows, after a close to linear time preprocessing of the input dataset, to find an approximately nearest neighbor of any fixed query in sublinear time in the dataset size. The resulting data structure is randomized and succeeds with high probability for every fixed query. In many modern applications of nearest neighbor search the queries are chosen adaptively. In this paper, we study the robustness of the locality-sensitive hashing to adaptive queries in Hamming space. We present a simple adversary that can, under mild assumptions on the initial point set, provably find a query to the approximate near neighbor search data structure that the data structure fails on. Crucially, our adaptive algorithm finds the hard query exponentially faster than random sampling.
@@ 1 Introduction
Nearest neighbor search is one of the most fundamental data structure problems in machine learning. In particular, the development of efficient data structures for nearest neighbor search in high dimensions is a challenge and finding the exact nearest neighbor often requires a linear scan over the data [WSB98]. Therefore, people also started to investigate approximate nearest neighbor search. One important technique to design approximate nearest neighbor data structures is locality sensitive hashing (LSH) which has been introduced in the seminal paper by Indyk and Motwani [IM98]. LSH is widely used in many applications including music retrieval [RK08], video anomaly detection [${}^{+}$16], and clustering [KIW07] (see also the survey [${}^{+}$21] for further examples).
LSH is a randomized data structure and as such its performance guarantees depend on its internal randomness. The standard analysis assumes an oblivious adversary, i.e. queries may not depend on ansers to earlier queries for the guarantees to apply. What are the consequences, if we would like to have the LSH performance guarantees when LSH is used as a data structure in an algorithm? It essentially says that for the guarantees to hold, the algorithm must be able to specify at a single point of time a set $Q$ of query points and it gets all answers to these query points at once. Then it can do any computation with the answers but it may not ask further queries. Such a restricted functionality significantly limits the strength of the guarantees provided by LSH as more often than not future queries depend on the result of earlier queries. Therefore, it would be great to better understand the LSH guarantees in the adaptive setting. Even though some work on LSH such as constructions without false negatives [Pag16, Ahl17] (see also Related Work for further discussions) are motivated by this question, there is still a lack of knowledge about how resilient standard LSH data structures are if we allow adaptive adversaries.
@@
Let us discuss this further on a simple example. We may use LSH for an approximate version of a $1$-nearest neighbor classifier (see [ABVT20] for a related study), that is, we store the training set in an LSH data structure and on a query we return the label of the (approximate) nearest point in the training set that is returned by LSH. In this setting LSH will satisfy its standard guarantees, when the choice of queries does not depend on previous answers, for example, if one thinks of the queries to originate iid. from a fixed query distribution. However, one could potentially design a carefully engineered sequence of queries to construct queries on which the algorithm returns an incorrect answer. This is particularly problematic if the algorithm is used for a sensitive task and relies on strong guarantees.
Our results.
Our main result is an algorithm for generating an adversarial query:
@@
We then evaluate our data structure experimentally in various settings to understand how the efficiency and effectiveness of our approach depends on the various parameters.
@@ 2 Related Work
Most of the work dedicated to Locality-Sensitive Hashing was done in the Monte-Carlo model, which we consider to be the standard model in this paper. That is, the goal is to construct a data structure with deterministically bounded query time that correctly answers the queries with high probability. Besides the [IM98] construction for Hamming space, locality-sensitive hash families have been constructed for several different metrics, including $\ell_{1}$ [DIIM04], the Euclidean metric [AI06], Jaccard similarity [Cha02], angular distance [${}^{+}$15] and more. Asymmetric LSH solutions, which allow to reduce query time at the expense of an increase in space, or reduce space at the expense of increasing query time, have received significant attention in the literature and are quite well understood by now [Pan06, Kap15, Chr17, ALRW17]. Distance sensitive hash families, where collision probabilities are not necessarily monotone in the distance between points being hashed, have also been constructed [ACPS18]. Data-dependent results, which adapt the locality-sensitive hash family to the dataset to achieve better exponents, are also known [AINR14, ALRW17].
Las-Vegas Locality Sensitive Hashing.
@@
In an attempt to avoid the problem of having false negatives, another line of work seeks to construct LSH instances that answer the queries deterministically correctly, but the running time is a random variable and guarantees are usually on its expectation. It was first shown by [Pag16] that such a scheme exists, however this scheme for general parameter settings requires more space than the standard LSH. A later work by [Ahl17] achieves the same space-time trade-off as in [IM98]. [Wei22] further improves on it for $\ell_{p}$, and offers a data-dependent construction following [AI06]. Although Las Vegas constructions avoid the problem of having false negatives, their guarantees still hold only against an oblivious adversary. It is entirely possible that an adaptive adversary can find a sequence of queries with an average response time much larger than what is guaranteed against an oblivious adversary. In any case, investigation of such adversaries lies outside of the scope of this paper. In [KOR98] a different approach to approximate nearest neighbor search was considered. Their data structure is false-negative free with high probability and supports $(1+\epsilon)$-approximate nearest neighbor queries with logarithmic query time while using polynomial space.
Robustness to adaptive adversaries.
@@
One of the main motivations of this paper is to study the behavior of a classic data structure solving against an adaptive adversary. One of the most studied class algorithms that is relevant to our paper is linear sketches. Some of the notable early works here include [HW13], which proposed an efficient attack on the linear sketch for norm estimation, and [NNW12], which proposed deterministic sketch constructions for a range of problems that work for all inputs and provide lower bounds on their size. More recently, [CN20] gave a robust construction for vector norm estimation, while an adaptive attack on CountSketch was proposed by [CNSS23, ${}^{+}$22], with the latter also proposing a robustifycation of this sketch.
@@
Another fruitful line of work covers the use of differential privacy to make existing algorithms more robust against adaptive adversaries [${}^{+}$22, WZ21, ${}^{+}$22, BJWY22, ACSS23]. The main idea is to use differentially private mechanisms to hide the random bits of the algorithm from the adversary. This approach turned out to be applicable to a wide range of estimation problems.
@@ 3 Preliminaries
Basics of Locality Sensitive Hashing.
Locality sensitive hashing is a technique to derive randomized data structures for near neighbor search. This can be formalized as follows. A (randomized) data structure for the $(c,cr)$-approximate near neighbor problem in a metric space $(X,\operatorname*{dist})$ preprocesses a point set $P\subset X$ in such a way that for a query point $q\in X$ with probability at least $1-\delta$ it (a) returns $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, if there is $p^{\prime}\in P$ with $\operatorname*{dist}(p^{\prime},q)\leq r$, (b) returns $\bot$, if there is no point $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, and (c) returns either $\bot$ or $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, otherwise. Here $\delta$ is the failure probability of the data structure. In this paper, we will consider the Hamming space, i.e. $X=\{0,1\}^{d}$ and $\operatorname*{dist}$ is the Hamming distance and in the following discussion we focus on this setting.
@@
For a $d$-dimensional Hamming space $\{0,1\}^{d}$, the locality-sensitive hash family has a very simple form [IM98]. Let =$\mathcal{H}=\{h^{(i)},i\in[d]:\forall p\in\{0,1\}^{d},h^{(i)}(p)=p_{i}\}$ be the elementary hash function family. That is, it consists of hash functions $h^{(i)}:\{0,1\}^{d}\to\{0,1\}$, which project points from $\{0,1\}$ to their $i$-th bit. It is known that
Fact 3.2 ([IM98]).
@@
The proof of [IM98] essentially follows by choosing the parameters $k$ and $L$ appropriately, which is done as follows. We set $p_{1}=1-r/d$ and $p_{2}=1-cr/d$ such that $\mathcal{H}$ is $(r,cr,p_{1},p_{2})$-sensitive. We define $\rho=\frac{\log 1/p_{1}}{\log 1/p_{2}}$ and $\ell=n^{\rho}$. For an input set $P\subseteq\{0,1\}^{d}$ of $n$ points we set $k=\lceil\log_{1/p_{2}}n\rceil$. For this setting of parameter, we get the guarantees provided in the above Theorem. In order to get a smaller failure probability we set $L=\lceil\lambda\ell\rceil$, which results in a failure probability of $1-\exp({-\Omega(\lambda)})$.
@@
Defining the adversary.
The goal of this paper is to study how one can force LSH to return a false negative by doing adaptive queries. We take the role as an adversary and our goal is design an adversarial algorithm that has access to an LSH instance and quickly forces it to return a false negative. In particular, we provide guarantees on the number of queries required until LSH makes a false negative.
@@ 4 Adversarial Algorithm
Algorithm outline.
The idea of the algorithm is to query a random point $q$ at distance $r-t$ such that the expected number of hash functions hashing $q$ to the same bucket as $z$ is at most $t/2$. We show that this is true when $t=2e^{2}(\lambda+1)$, which is the value we use throughout this section. This implies that with a constant probability $|\textsc{Coll}(q,z)|\leq t$. The algorithm now does $t$ iterations and in each of them it moves point $q$ away from $z$ by flipping one bit in it in such a way that the size of the set $\textsc{Coll}(q,z)$ also decreases by at least $1$.
To find this bit, the algorithm flips bits in $q$ at $cr-\operatorname*{dist}(q,z)$ randomly chosen positions among those where $q$ is equal to $z$ and obtains a point $\tilde{q}$. $\tilde{q}$ is located at distance $cr$ from $z$, so with a high probability $\textsc{QueryLSH}(\tilde{q})=\bot$. Now consider a path between $q$ and $\tilde{q}$, i.e. a sequence of points starting with $q$ and ending with $\tilde{q}$, where each two consecutive points differ in one bit and where each point is further away from $z$ than its predecessor. Because $\tilde{q}$ doesn’t hash with $z$ but $q$ does, there are two consecutive points $q^{\prime}$, $q^{\prime\prime}$ on this path with the same property. Consider the bit where they differ. Because $\textsc{Coll}(q^{\prime},z)\neq\emptyset$ and $\textsc{Coll}(q^{\prime\prime},z)=\emptyset$, this bit must lie in the support of all functions in $\textsc{Coll}(q^{\prime},z)$222 for an elementary hash function $h^{(i)}\in\mathcal{H}$ its support is defined as the set $\{i\}$. For a hash function $g\in G$, $g=(h_{1},\ldots,h_{k})$ its support is the union of supports of $h_{1},\ldots,h_{k}$.. Therefore, by flipping it in $q$ the algorithm removes these functions from $\textsc{Coll}(q,z)$. It is easy to see that at the end we get a point $q$ at a distance at most $r$ from $z$ such that $\textsc{Coll}(q,z)=\emptyset$, so $\textsc{QueryLSH}(q)=\bot$ and $q$ is a false negative.
4.1 Key lemmas
In this section we prove Lemmas 4.3 and 4.4, which are the key results underpinning the correctness of our algorithm. Lemma 4.3 essentially shows that a point $q$ sampled at distance $r-t$ collides with $z$ in at most $O(t)$ hashings with constant probability. Lemma 4.4 that says that if we move this point randomly and monotonically away from $z$ to distance $cr$, it will not hash at all with $z$.
@@ 4.1 Key lemmas
Let $n>e$, $8e^{2}(\lambda+1)\ln n\leq cr$, $28\ln^{3}n/c^{2}\leq r\leq d/(14\ln n)$ and $r\leq d/(5c)$. Let the hash functions $G$ be such that the size of the support of each is at least $k-7\ln n\cdot\max\{1,\frac{k^{2}}{2d}\}$. Let $t=2e^{2}(\lambda+1)$. Let $q$ be a point at distance $r-t$ from $z$ chosen uniformly at random. The expected size of $\textsc{Coll}(q,z)$ is at most $e^{2}(\lambda+1)$, where the expectation is taken with respect to the randomness of $q$.
@@
$\operatorname*{\mathbb{E}}[|\textsc{Coll}(q,z)|]=\sum_{g\in G}\operatorname*{\mathbb{E}}[X_{g}]=\sum_{g\in G}\Pr[g(q)=q(z)],$
where $X_{g}$ is the indicator random variable for the event $g(q)=g(z)$ and the randomness is over the choice of $q$ as in the lemma.
In the following we consider an arbitrary $g\in G$ and derive an upper bound for the collision probability of $q$ and $z$. Indeed, let $x=7\ln n\cdot\max\{1,\frac{k^{2}}{2d}\}$. We first observe (see Section A.1 for more details) that $\Pr[g(q)=g(z)]\leq\left(1-\frac{k-x}{d}\right)^{r-t}.$ The basic argument is that sampling a point at distance $r$ from $z$ is not worse than just picking $r$ bits to flip in $z$ independently at random, and then bounding the probability that none of those chosen bits lie in the support of $g$.
Next we can use this upper bound to argue that
$\sum_{g\in G}\Pr[g(q)=g(z)]\leq L\cdot\left(1-\frac{k-x}{d}\right)^{r-t}.$
Using $L=\lceil\lambda\ell\rceil\leq(\lambda+1)\ell$ and $\ell=n^{\rho}=\left(1-\frac{r}{d}\right)^{-\log_{1/p_{2}}n}\leq\left(1-\frac{r}{d}\right)^{-k}$ we obtain
$\begin{split}\operatorname*{\mathbb{E}}_{q}\left[|\textsc{Coll}(q,z)|\right]&\leq(\lambda+1)\left(1-\frac{r}{d}\right)^{-k}\left(1-\frac{k-x}{d}\right)^{r-t}\\ \end{split}$
It remains to show that the r.h.s. is at most $e^{2}(\lambda+1)$. Because $x$ and $t$ are considerably smaller than $k$ and $r$ respectively, this essentially boils down to show that $\left(1-\frac{r}{d}\right)^{-k}\cdot\left(1-\frac{k}{d}\right)^{r}\leq O(1).$ This is simple to bound using the Taylor expansion of both terms (see Section A.1 for detailed calculations). ∎
@@
Fix $g\in\textsc{Coll}(q,z)$. First notice that $q^{\prime}$ defined in lemma statement can be generated as follows. One first selects a uniformly random subset $S$ of $cr-\operatorname*{dist}(q,z)$ coordinates in $I=\{i\in[d]:q_{i}=z_{i}\}$. Then one lets $q^{\prime}_{i}=q_{i}\oplus 1$ (negation of $q_{i}$) for $i\in S$ and $q^{\prime}_{i}=q_{i}$ for $i\in[d]\setminus S$. We denote this distribution by $\mathcal{D}(q)$. Note that because $g(q)=g(z)$ (since $g\in\textsc{Coll}(q,z)$ by assumption), the support of $g$ is a subset of $I$. We will show the following
$\Pr_{q^{\prime}\sim\mathcal{D}(q)}[g(q^{\prime})=g(z)]\leq\left(1-\frac{k}{2d}\right)^{(c-1)r}.$ (1)
In order to establish the above, we view the generation of $q^{\prime}$ as a two step process. First we uniformly sample $cr-\operatorname*{dist}(q,z)\geq cr-r$ bits from $I$ with replacement into a set $B$ (removing duplicates) and flip them in $q$ – denote the result by $\tilde{q}$. Next, we choose $cr-\operatorname*{dist}(\tilde{q},z)$ more bits from $I\setminus B$ and flip them in $\tilde{q}$ to get $q^{\prime}$. It again holds that $g(\tilde{q})\neq g(z)$ implies $g(q^{\prime})\neq g(z)$, so $\Pr_{q^{\prime}}(g(q^{\prime})\neq g(z))\geq\Pr_{\tilde{q}}(g(\tilde{q})\neq g(z))$.
The probability that each bit chosen in the first step lies in support of $g$ is at least $\frac{k/2}{d-\operatorname*{dist}(q,r)}\geq\frac{k}{2d}$. Hence, the probability that none of them lie in the support, i.e. $\Pr_{\tilde{q}}(g(\tilde{q})=g(z))$, is at most $(1-\frac{k}{2d})^{cr-r}$. This implies (1).
By Lemma 4.1, $k\geq 0.5\frac{d}{cr}\ln$. By applying A.1 to $(1-\frac{k}{2d})^{cr-r}$, we obtain:
$\Pr[g(q^{\prime})=g(z)]\leq\left(1-\frac{k}{2d}\right)^{(c-1)r}\leq n^{-\frac{c-1}{4c}}.$ (2)
Now consider two cases. First, if $c\geq 2$, then the rhs of (2) satisfies $n^{-\frac{c-1}{4c}}\leq n^{-1/4}\leq\frac{1}{(8e^{2}(\lambda+1))^{2}}$ since $8e^{2}(\lambda+1)\leq n^{1/8}$ by assumption of the lemma. Otherwise, using that $c\geq 1+\frac{80+16\ln(\lambda+1)}{\ln n}$ by the assumption of the lemma, we get $\frac{c-1}{4c}\geq\frac{c-1}{8}\geq\frac{10+2\ln(\lambda+1)}{\ln n}\geq\ln(8e^{2}(\lambda+1))/\ln n$.
Thus, in both cases one has $n^{-\frac{c-1}{4c}}\leq\frac{1}{(8e^{2}(\lambda+1))^{2}}$. Finally, since $|\textsc{Coll}(q,z)|\leq 2e^{2}(\lambda+1)$ by the assumption of the lemma, by union bound across all $g\in\textsc{Coll}(q,z)$, $\Pr[\textsc{Coll}(q^{\prime},z)\neq\emptyset]\leq 1/(32e^{2}(\lambda+1))$, as required. ∎

### v3

@@ Abstract
Locality-sensitive hashing [22] is a classical data structure for approximate nearest neighbor search. It allows, after a close to linear time preprocessing of the input dataset, to find an approximately nearest neighbor of any fixed query in sublinear time in the dataset size. The resulting data structure is randomized and succeeds with high probability for every fixed query. In many modern applications of nearest neighbor search the queries are chosen adaptively. In this paper, we study the robustness of the locality-sensitive hashing to adaptive queries in Hamming space. We present a simple adversary that can, under mild assumptions on the initial point set, provably find a query to the approximate near neighbor search data structure that the data structure fails on. Crucially, our adaptive algorithm finds the hard query exponentially faster than random sampling.
@@ 1 Introduction
Nearest neighbor search is one of the most fundamental data structure problems in machine learning. In particular, the development of efficient data structures for nearest neighbor search in high dimensions is a challenge and finding the exact nearest neighbor often requires a linear scan over the data [35]. Therefore, people also started to investigate approximate nearest neighbor search. One important technique to design approximate nearest neighbor data structures is locality sensitive hashing (LSH) which has been introduced in the seminal paper by Indyk and Motwani [22]. LSH is widely used in many applications including music retrieval [33], video anomaly detection [39], and clustering [26] (see also the survey [23] for further examples).
LSH is a randomized data structure and as such its performance guarantees depend on its internal randomness. The standard analysis assumes an oblivious adversary, i.e. queries may not depend on answers to earlier queries for the guarantees to apply. What are the consequences, if we would like to have the LSH performance guarantees when LSH is used as a data structure in an algorithm? It essentially says that for the guarantees to hold, the algorithm must be able to specify at a single point of time a set $Q$ of query points and it gets all answers to these query points at once. Then it can do any computation with the answers but it may not ask further queries. Such a restricted functionality significantly limits the strength of the guarantees provided by LSH as more often than not future queries depend on the result of earlier queries. Therefore, it would be great to better understand the LSH guarantees in the adaptive setting. Even though some work on LSH such as constructions without false negatives [31, 1] (see also Related Work for further discussions) are motivated by this question, there is still a lack of knowledge about how resilient standard LSH data structures are if we allow adaptive adversaries.
@@
Let us discuss this further on a simple example. We may use LSH for an approximate version of a $1$-nearest neighbor classifier (see [2] for a related study), that is, we store the training set in an LSH data structure and on a query we return the label of the (approximate) nearest point in the training set that is returned by LSH. In this setting LSH will satisfy its standard guarantees when the choice of queries does not depend on previous answers. This happens, for example, when queries are independently sampled from a fixed query distribution. However, one could potentially design a carefully engineered sequence of queries to construct queries on which the algorithm returns an incorrect answer. This is particularly problematic if the algorithm is used for a sensitive task and relies on strong guarantees.
Our results. Our main result is an algorithm for generating an adversarial query:
@@
Note that to find such a point through random sampling, one would have to make $\Omega(1/\delta)$ queries in expectation Theorem 3.3, which is exponentialy slower in $\delta$ than the proposed adversary. We then evaluate our data structure experimentally in various settings to understand how the efficiency and effectiveness of our approach depends on the various parameters.
@@ 2 Related Work
Monte-Carlo Locality-Sensitive Hashing. Most of the work dedicated to Locality-Sensitive Hashing was done in the Monte-Carlo model, which we consider to be the standard model in this paper. That is, the goal is to construct a data structure with deterministically bounded query time that correctly answers the queries with high probability. Constructions of locality-sensitive hash families are known for several different metrics, including Hamming distance [22, 7], $\ell_{1}$ [19], the Euclidean metric [3], Jaccard similarity [13, 18], angular distance [4, 25, 29] and more. Asymmetric LSH solutions, which allow to reduce query time at the expense of an increase in space, or reduce space at the expense of increasing query time, have received significant attention in the literature and are quite well understood by now [32, 24, 15, 6]. Distance sensitive hash families, where collision probabilities are not necessarily monotone in the distance between points being hashed, have also been constructed [9]. Data-dependent results, which adapt the locality-sensitive hash family to the dataset to achieve better exponents, are also known [5, 6].
@@
Las-Vegas Locality Sensitive Hashing. In an attempt to avoid the problem of having false negatives, another line of work seeks to construct LSH instances that answer the queries deterministically correctly, but the running time is a random variable and guarantees are usually on its expectation. It was first shown by [31] that such a scheme exists, however this scheme for general parameter settings requires more space than the standard LSH. A later work by [1] achieves the same space-time trade-off as in [22]. [36] further improves on it for $\ell_{p}$, and offers a data-dependent construction following [3]. Although Las Vegas constructions avoid the problem of having false negatives, their guarantees still hold only against an oblivious adversary. It is entirely possible that an adaptive adversary can find a sequence of queries with an average response time much larger than what is guaranteed against an oblivious adversary. In any case, investigation of such adversaries lies outside of the scope of this paper.
@@
Robustness to adaptive adversaries. One of the main motivations of this paper is to study the behavior of a classic data structure solving against an adaptive adversary. One of the most studied class of algorithms that is relevant to our paper is linear sketches. Some of the notable early works here include [20], which proposed an efficient attack on the linear sketch for norm estimation, and [30], which proposed deterministic sketch constructions for a range of problems that work for all inputs and provide lower bounds on their size. In [27] an adversarially robust approach to solving approximate nearest neighbor in Hamming space is proposed, which does not use LSH. It is improved upon by [14], which gives a robust construction for vector $\ell_{p}$-norm estimation, for $0<p\leq 2$. On the other side of the spectrum, an adaptive attack on CountSketch was proposed by [17, 16], with the latter also proposing a robustifycation of this sketch.
@@
Another fruitful line of work covers the use of differential privacy to make existing algorithms more robust against adaptive adversaries [10, 37, 21, 11, 8]. The main idea is to use differentially private mechanisms to hide the random bits of the algorithm from the adversary. This approach turned out to be applicable to a wide range of estimation problems.
@@ 3 Preliminaries
Basics of Locality Sensitive Hashing. Locality sensitive hashing is a technique to derive randomized data structures for near neighbor search. This can be formalized as follows. A (randomized) data structure for the $(c,cr)$-approximate near neighbor problem in a metric space $(X,\operatorname*{dist})$ preprocesses a point set $P\subset X$ in such a way that for a query point $q\in X$ with probability at least $1-\delta$ it (a) returns $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, if there is $p^{\prime}\in P$ with $\operatorname*{dist}(p^{\prime},q)\leq r$, (b) returns $\bot$, if there is no point $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, and (c) returns either $\bot$ or $p\in P$ with $\operatorname*{dist}(p,q)\leq cr$, otherwise. Here $\delta$ is the failure probability of the data structure. In this paper, we will consider the Hamming space, i.e. $X=\{0,1\}^{d}$ and $\operatorname*{dist}$ is the Hamming distance and in the following discussion we focus on this setting.
@@
For a $d$-dimensional Hamming space $\{0,1\}^{d}$, the locality-sensitive hash family has a very simple form [22]. Let =$\mathcal{H}=\{h^{(i)},i\in[d]:\forall p\in\{0,1\}^{d},h^{(i)}(p)=p_{i}\}$ be the elementary hash function family. That is, it consists of hash functions $h^{(i)}:\{0,1\}^{d}\to\{0,1\}$, which project points from $\{0,1\}$ to their $i$-th bit. It is known that
Fact 3.2 ([22]).
@@
The proof of [22] essentially follows by choosing the parameters $k$ and $L$ appropriately, which is done as follows. We set $p_{1}=1-r/d$ and $p_{2}=1-cr/d$ such that $\mathcal{H}$ is $(r,cr,p_{1},p_{2})$-sensitive. We define $\rho=\frac{\log 1/p_{1}}{\log 1/p_{2}}$ and $\ell=n^{\rho}$. For an input set $P\subseteq\{0,1\}^{d}$ of $n$ points we set $k=\lceil\log_{1/p_{2}}n\rceil$. For this setting of parameter, we get the guarantees provided in the above Theorem. In order to get a smaller failure probability we set $L=\lceil\lambda\ell\rceil$, which results in a failure probability of $1-\exp({-\Omega(\lambda)})$.
@@
Defining the adversary. The goal of this paper is to study how one can force LSH to return a false negative by doing adaptive queries. We take the role as an adversary and our goal is design an adversarial algorithm that has access to an LSH instance and quickly forces it to return a false negative. In particular, we provide guarantees on the number of queries required until LSH makes a false negative.
@@ 4 Adversarial Algorithm
Algorithm outline. The idea of the algorithm is to query a random point $q$ at distance $r-t$ such that the expected number of hash functions hashing $q$ to the same bucket as $z$ is at most $t/2$. We show that this is true when $t=2e^{2}(\lambda+1)$, which is the value we use throughout this section. This implies that with a constant probability $|\textsc{Coll}(q,z)|\leq t$. The algorithm now does $t$ iterations and in each of them it moves point $q$ away from $z$ by flipping one bit in it in such a way that the size of the set $\textsc{Coll}(q,z)$ also decreases by at least $1$.
To find this bit, the algorithm flips bits in $q$ at $cr-\operatorname*{dist}(q,z)$ randomly chosen positions among those where $q$ is equal to $z$ and obtains a point $\tilde{q}$. $\tilde{q}$ is located at distance $cr$ from $z$, so with a high probability $\textsc{QueryLSH}(\tilde{q})=\bot$. Now consider a path between $q$ and $\tilde{q}$, i.e. a sequence of points starting with $q$ and ending with $\tilde{q}$, where each two consecutive points differ in one bit and where each point is further away from $z$ than its predecessor. Because $\tilde{q}$ doesn’t hash with $z$ but $q$ does, there are two consecutive points $q^{\prime}$, $q^{\prime\prime}$ on this path with the same property. Consider the bit where they differ. Because $\textsc{Coll}(q^{\prime},z)\neq\emptyset$ and $\textsc{Coll}(q^{\prime\prime},z)=\emptyset$, this bit must lie in the support of all functions in $\textsc{Coll}(q^{\prime},z)$.111 for an elementary hash function $h^{(i)}\in\mathcal{H}$ its support is defined as the set $\{i\}$. For a hash function $g\in G$, $g=(h_{1},\ldots,h_{k})$ its support is the union of supports of $h_{1},\ldots,h_{k}$. Therefore, by flipping it in $q$ the algorithm removes these functions from $\textsc{Coll}(q,z)$. It is easy to see that at the end we get a point $q$ at a distance at most $r$ from $z$ such that $\textsc{Coll}(q,z)=\emptyset$, so $\textsc{QueryLSH}(q)=\bot$ and $q$ is a false negative.
In the following we prove Lemmas 4.3 and 4.4, which are the key results underpinning the correctness of our algorithm. Lemma 4.3 essentially shows that a point $q$ sampled at distance $r-t$ collides with $z$ in at most $O(t)$ hashings with constant probability. Lemma 4.4 that says that if we move this point randomly and monotonically away from $z$ to distance $cr$, it will not hash at all with $z$.
@@
Let $n>e$, $8e^{2}(\lambda+1)\ln n\leq cr$, $28\ln^{3}n/c^{2}\leq r\leq d/(14\ln n)$, $r\leq d/(5c)$ and $c<\ln n$. Let the hash functions $G$ be such that the size of the support of each is at least $k-7\ln n\cdot\max\{1,\frac{k^{2}}{2d}\}$. Let $t=2e^{2}(\lambda+1)$. Let $q$ be a point at distance $r-t$ from $z$ chosen uniformly at random. The expected size of $\textsc{Coll}(q,z)$ is at most $e^{2}(\lambda+1)$, where the expectation is taken with respect to the randomness of $q$.
@@
We fix an arbitrary set $G$ of $L$ hash functions that satisfy the condition of the lemma. By linearity of expectation, $\operatorname*{\mathbb{E}}[|\textsc{Coll}(q,z)|]=\sum_{g\in G}\operatorname*{\mathbb{E}}[X_{g}]=\sum_{g\in G}\Pr[g(q)=q(z)],$ where $X_{g}$ is the indicator random variable for the event $g(q)=g(z)$ and the randomness is over the choice of $q$ as in the lemma. In the following we consider an arbitrary $g\in G$ and derive an upper bound for the collision probability of $q$ and $z$. Indeed, let $x=7\ln n\cdot\max\{1,\frac{k^{2}}{2d}\}$. We first observe (see Section A.1 for more details) that $\Pr[g(q)=g(z)]\leq\left(1-\frac{k-x}{d}\right)^{r-t}.$ The basic argument is that sampling a point at distance $r$ from $z$ is not worse than just picking $r$ bits to flip in $z$ independently at random, and then bounding the probability that none of those chosen bits lie in the support of $g$.
Next we can use this upper bound to argue that $\sum_{g\in G}\Pr[g(q)=g(z)]\leq L\cdot\left(1-\frac{k-x}{d}\right)^{r-t}.$ Using $L=\lceil\lambda\ell\rceil\leq(\lambda+1)\ell$ and $\ell=n^{\rho}=\left(1-\frac{r}{d}\right)^{-\log_{1/p_{2}}n}\leq\left(1-\frac{r}{d}\right)^{-k}$ we obtain $\operatorname*{\mathbb{E}}_{q}\left[|\textsc{Coll}(q,z)|\right]\leq(\lambda+1)\left(1-\frac{r}{d}\right)^{-k}\left(1-\frac{k-x}{d}\right)^{r-t}.$ It remains to show that the r.h.s. is at most $e^{2}(\lambda+1)$. Because $x$ and $t$ are considerably smaller than $k$ and $r$ respectively, this essentially boils down to show that $\left(1-\frac{r}{d}\right)^{-k}\cdot\left(1-\frac{k}{d}\right)^{r}\leq O(1).$ This is simple to bound using the Taylor expansion of both terms (see Section A.1 for detailed calculations). ∎
@@
Fix $g\in\textsc{Coll}(q,z)$. First notice that $q^{\prime}$ defined in lemma statement can be generated as follows. One first selects a uniformly random subset $S$ of $cr-\operatorname*{dist}(q,z)$ coordinates in $I=\{i\in[d]:q_{i}=z_{i}\}$. Then one lets $q^{\prime}_{i}=q_{i}\oplus 1$ (negation of $q_{i}$) for $i\in S$ and $q^{\prime}_{i}=q_{i}$ for $i\in[d]\setminus S$. We denote this distribution by $\mathcal{D}(q)$. Note that because $g(q)=g(z)$ (since $g\in\textsc{Coll}(q,z)$ by assumption), the support of $g$ is a subset of $I$. We will show that $\Pr_{q^{\prime}\sim\mathcal{D}(q)}[g(q^{\prime})=g(z)]\leq\left(1-\frac{k}{2d}\right)^{(c-1)r}.$ In order to establish the above, we view the generation of $q^{\prime}$ as a two step process. First we uniformly sample $cr-\operatorname*{dist}(q,z)\geq cr-r$ bits from $I$ with replacement into a set $B$ (removing duplicates) and flip them in $q$ – denote the result by $\tilde{q}$. Next, we choose $cr-\operatorname*{dist}(\tilde{q},z)$ more bits from $I\setminus B$ and flip them in $\tilde{q}$ to get $q^{\prime}$. It again holds that $g(\tilde{q})\neq g(z)$ implies $g(q^{\prime})\neq g(z)$, so $\Pr_{q^{\prime}}(g(q^{\prime})\neq g(z))\geq\Pr_{\tilde{q}}(g(\tilde{q})\neq g(z))$.
The probability that each bit chosen in the first step lies in support of $g$ is at least $\frac{k/2}{d-\operatorname*{dist}(q,r)}\geq\frac{k}{2d}$. Hence, the probability that none of them lie in the support, i.e. $\Pr_{\tilde{q}}(g(\tilde{q})=g(z))$, is at most $(1-\frac{k}{2d})^{cr-r}$. This establishes the inequality.
By Lemma 4.1, $k\geq 0.5\frac{d}{cr}\ln$. By applying A.1 to $(1-\frac{k}{2d})^{cr-r}$, we obtain: $\Pr[g(q^{\prime})=g(z)]\leq\left(1-\frac{k}{2d}\right)^{(c-1)r}\leq n^{-\frac{c-1}{4c}}.$ Now consider two cases. First, if $c\geq 2$, then the rhs of the above inequality satisfies $n^{-\frac{c-1}{4c}}\leq n^{-1/4}\leq\frac{1}{(8e^{2}(\lambda+1))^{2}}$ since $8e^{2}(\lambda+1)\leq n^{1/8}$ by assumption of the lemma. Otherwise, using that $c\geq 1+\frac{80+16\ln(\lambda+1)}{\ln n}$ by the assumption of the lemma, we get $\frac{c-1}{4c}\geq\frac{c-1}{8}\geq\frac{10+2\ln(\lambda+1)}{\ln n}\geq\ln(8e^{2}(\lambda+1))/\ln n$. Thus, in both cases one has $n^{-\frac{c-1}{4c}}\leq\frac{1}{(8e^{2}(\lambda+1))^{2}}$. Finally, since $|\textsc{Coll}(q,z)|\leq 2e^{2}(\lambda+1)$ by the assumption of the lemma, by union bound across all $g\in\textsc{Coll}(q,z)$, $\Pr[\textsc{Coll}(q^{\prime},z)\neq\emptyset]\leq 1/(32e^{2}(\lambda+1))$, as required. ∎